OpenStudy (miasmith):

Solve systems using matrices practice, please help, will post questions below

OpenStudy (miasmith):

1. write a matrix to represent the system 2a-3b=6 a+b=2

OpenStudy (wolf1728):

I call these determinants. It's the same thing right?

OpenStudy (sooobored):

no

OpenStudy (wolf1728):

Let's suppose we have 2 equations: ax + by = e cx + dy = f the denominator determinant is a b c d evaluated as a*d - c*b

OpenStudy (miasmith):

these are the options A: 2 -3 6 1 1 2 B: 2 1 -3 1 6 2 C: a b 6 a b 2 D: 2a -3b 6 a b 2 all in brackets

OpenStudy (sooobored):

yes that is the determinant, the question is just asking to express it in terms of matrices or in the form of Ax=B where \[A=\left[\begin{matrix}2 & -3 \\ 1 & 1\end{matrix}\right]\]\[x=\left(\begin{matrix}a \\ b\end{matrix}\right)\] \[B=\left(\begin{matrix}6 \\ 2\end{matrix}\right)\]

OpenStudy (wolf1728):

2a-3b=6 a+b=2 2 -3 6 1 1 2 Looks like choice A to me.

OpenStudy (wolf1728):

a=2 b=-3 e=6 c=1 d=1 f=2 denominator = a*d -c*b = 2 -1*-3 = 5 x = 6 -3 2 1 =6 --6 = 12 divide by denominator 5 x = 12/5 = 2.4